# Symbolic Powers via Extension

### v1

@@ 1. Introduction
Let $A$ and $B$ be Noetherian rings and $\phi:A\to B$ be a flat morphism. Assume that $I\subset A$ is a proper ideal such that $\operatorname{Ass}_{B}(IB)=\{{{\mathfrak{p}}}B~{}|~{}{{\mathfrak{p}}}\in\operatorname{Ass}_{A}(I)\}$, and every associated prime of $I$ is a contraction of some ideal in $B$. Then $(IB)^{(m)}=I^{(m)}B$.
@@
One of the central topic of interest in the study of symbolic powers is to find $i$ and $j$ for which $I^{(i)}\subset I^{j}$. In [ELS01, HH02, MS18], it has been proved that if $I$ is a homogeneous ideal of bigheight $h$ in a polynomial ring then $I^{(rh)}\subset I^{r}$ for all $r\geqslant 1$.
In the other hand, one can investigate non-containment between the symbolic and ordinary powers of ideals. The resurgence, defined by Bocci and Harbourne [BH10], and the asymptotic resurgence, defined by Guardo, Harbourne, and Van Tuyl [GHVT13] measure such non-containment.
If $I$ is a homogeneous ideal in a finitely generated ${\mathbbm{k}}$-algebra domain $A$, then the resurgence number, $\rho(I)$ and the asymptotic resurgence, $\rho_{a}(I)$ are defined as follows:
@@
If $I\subset{\mathbbm{k}}[\underline{x}]$ and $J\subset{\mathbbm{k}}[\underline{y}]$ are ideals in polynomial rings ${\mathbbm{k}}[\underline{x}]$ and ${\mathbbm{k}}[\underline{y}]$ then the resurgence and the asymptotic resurgence numbers of $I+J\subset{\mathbbm{k}}[\underline{x},\underline{y}]$ attracted huge attention.
Motivated by 1.1 we study the following question.
@@
Let $A$ and $B$ be noetherian rings, $I$ be an ideal of $A$, and $J$ be an ideal of $B$. What can be said about $\rho(I+J)$ and $\rho_{a}(I+J)$, where $I+J=IR+JR$, where $R=A\otimes B$?
The resurgence and the asymptotic resurgence of the sum of ideals $I\subset{\mathbbm{k}}[\underline{x}]$ and $J\subset{\mathbbm{k}}[\underline{y}]$, where ${\mathbbm{k}}[\underline{x}]$ and ${\mathbbm{k}}[\underline{y}]$ are polynomial rings, are studied in [BHJT21, KNT24]. As an application of Theorem 2.5 and Corollary 3.4, we extend those results ideals in standard graded finitely algebra domains over algebraically closed fields.
@@
Let $I\subset A$ and $J\subset B$ be homogeneous ideals in finitely generated ${\mathbbm{k}}$-algebra domains $A$ and $B$, where ${\mathbbm{k}}$ is algebraically closed. Set $I+J=IR+JR$, where $R=A\otimes B$. Then the following hold .
@@ Acknowledgments.
The authors are thankful to Tài Huy Hà for suggesting the problem and valuable comments. This is a part of the second author’s Ph.D. thesis.
@@ 2. Symbolic Powers via flat extension
Let $\phi:A\to B$ be a flat homomorphism. If $I\subset A$ is a ideal then we have $\operatorname{Ass}(IB)=\cup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\operatorname{Ass}({\mathfrak{p}}B)$. Furthermore, if $\operatorname{Ass}(IB)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}(I)\}$ then we have $\operatorname{Ass}(I^{(n)}B)\subset\operatorname{Ass}(IB)$ for all $n\geqslant 1$.
@@
$\operatorname{Ass}(IB)=\operatorname{Ass}_{B}(A/I\otimes_{A}B)=\bigcup_{\mathfrak{p}\in\operatorname{Ass}_{A}(A/I)}\operatorname{Ass}_{B}(B/\mathfrak{p}B)=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\operatorname{Ass}_{B}({\mathfrak{p}}B).$
Note that $\operatorname{Ass}(I^{(n)})\subset\operatorname{Ass}(I)$ for all $n\geqslant 1$. Since $\operatorname{Ass}(IB)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}(I)\}$ then using Theorem 2.3 we get
$\operatorname{Ass}(I^{(n)}B)=\bigcup_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}\left(I^{(n)}\right)}\operatorname{Ass}({\mathfrak{p}}^{\prime}B)=\left\{{\mathfrak{p}}^{\prime}B~{}|~{}{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I^{(n)})\right\}\subset\operatorname{Ass}(IB),\text{ as }\operatorname{Ass}\left(I^{(n)}\right)\subset\operatorname{Ass}(I).$
@@
Let $A$ and $B$ be Noetherian rings and $\phi:A\to B$ be a flat morphism. Assume that $I\subset A$ is a proper ideal such that $\operatorname{Ass}_{B}(IB)=\{{{\mathfrak{p}}}B~{}|~{}{{\mathfrak{p}}}\in\operatorname{Ass}_{A}(I)\}$ and every associated prime of $I$ is a contraction of some ideal in $B$ Then $(IB)^{(n)}=I^{(n)}B$, for all $n\geqslant 1$.
@@
Using the definition of symbolic powers of an ideal and assumption, we get
@@
Now by Lemma 2.4, we have $\operatorname{Ass}\left(I^{(n)}B\right)\subset\operatorname{Ass}(IB)$. Thus whenever $s\notin W$ then $s\not\in{\mathfrak{q}}$, where ${\mathfrak{q}}$ is an associated prime of $I^{(n)}B$. Thus by [AM94, Proposition 4.4(iii)] we get
@@ 3. Resurgence and Prime extension
The following is from [Wal18, Proposition 2.1].
Proposition 3.1.
Let $A$ and $B$ be Noetherian rings and $\phi:A\to B$ be a faithfully flat morphism. If $\operatorname{Ass}(IB)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}(I)\}$, then $\rho(I)=\rho(IB)$.
Now we discuss some specific classes of morphisms which together with flatness gives faithfully flat. We start with the following definition.
@@
Since $\phi$ is prime extension then by Remark 3.3, it is faithfully flat. Thus using Theorem 2.5 we get that
@@ 4. Resurgence of sum of ideals
In this section we prove similar results for finitely generated ${\mathbbm{k}}$-algebra domains, where ${\mathbbm{k}}$ is an algebraically closed field. One of the key tools in [BHJT21, Theorem 2.6] and [KNT24, Theorem 2.3] is the binomial expansion of symbolic powers of ideals, which holds for ideals in ${\mathbbm{k}}$-algebras. The other tool is $\rho(IR)=\rho(I)$, this holds in the polynomial ring set up. To generalize [BHJT21, Theorem 2.6] and [KNT24, Theorem 2.3], we need that $\rho(IR)=\rho(I)$ in general domains, which is not known.
In Lemma 4.1, we prove that if $A$ and $B$ are finitely generated algebra domains over an algebraically closed field ${\mathbbm{k}}$, and $R=A\otimes B$, then $\rho(IR)=\rho(I)$, and $\rho(JR)=\rho(J)$, where $I\subset A,J\subset B$ are ideals.
@@
Now $\left(A/\mathfrak{p}\right)\otimes_{k}B$ is an integral domain by [Iit82, Lemma 1.54]. It is a well-known fact that an ideal is prime if only if the quotient ring is an integral domain, so $\mathfrak{p}R$ is prime, and thus, $\phi$ has prime extension property. The last equality follows from Corollary 3.4. ∎
@@
To see why we need the conditions on $A$ and $B$, we will need some examples and theorems. Now [Vám78, Theorem 11] implies that, if $A$ (as a field extension over ${\mathbbm{k}}$) is not finitely generated as a ${\mathbbm{k}}$-algebra, then $A\otimes_{\mathbbm{k}}A$ is not a Noetherian ring. For an example, $\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}$ is not a Notherian ring. Therefore, the condition being finitely generated is necessary. We also need $A$ and $B$ to be integral domains as ${\mathbbm{k}}$-algebras where ${\mathbbm{k}}$ is algebraically closed because of [Iit82, Lemma 1.54], this will make $A\otimes_{\mathbbm{k}}B$ to be an integral domain. It’s a famous example that $\mathbb{C}\otimes_{\mathbb{R}}\mathbb{C}$ is not an integral domain.

### v2

@@ 1. Introduction
Let $A$ and $B$ be Noetherian rings and $\phi:A\to B$ be a flat morphism. Assume that $I\subset A$ is a proper ideal such that $\operatorname{Ass}^{*}(IB)=\{{\mathfrak{p}}B~{}|~{}\text{ for all }{\mathfrak{p}}\in\operatorname{Ass}^{*}(I)\}$, and every associated prime of $I$ is a contraction of some ideal in $B$. Then $(IB)^{(n)}=I^{(n)}B$, for all $n\geqslant 1$.
@@
One of the central topic of interest in the study of symbolic powers is to find $i$ and $j$ for which $I^{(i)}\subset I^{j}$. In [ELS01, HH02, MS18], it has been proved that if $I$ is a homogeneous ideal of bigheight $h$ in a polynomial ring then $I^{(rh)}\subset I^{r}$ for all $r\geqslant 1$. In the other hand, one can investigate non-containment between the symbolic and ordinary powers of ideals. The resurgence, defined by Bocci and Harbourne [BH10], and the asymptotic resurgence, defined by Guardo, Harbourne, and Van Tuyl [GHVT13] measure such non-containment.
If $I$ is a homogeneous ideal in a finitely generated regular ${\mathbbm{k}}$-algebra domain $A$, where ${\mathbbm{k}}$ is perfect then the resurgence number, $\rho(I)$ and the asymptotic resurgence, $\rho_{a}(I)$ are defined as follows:
@@
If $I\subset{\mathbbm{k}}[\underline{x}]$ and $J\subset{\mathbbm{k}}[\underline{y}]$ are ideals in polynomial rings ${\mathbbm{k}}[\underline{x}]$ and ${\mathbbm{k}}[\underline{y}]$ then the resurgence and the asymptotic resurgence numbers of $I+J\subset{\mathbbm{k}}[\underline{x},\underline{y}]$ attracted huge attention. Motivated by 1.1 we study the following question.
@@
Let $A$ and $B$ be regular noetherian rings, $I$ be an ideal of $A$, and $J$ be an ideal of $B$. What can be said about $\rho(I+J)$ and $\rho_{a}(I+J)$, where $I+J=IR+JR$, where $R=A\otimes B$?
The resurgence and the asymptotic resurgence of the sum of ideals $I\subset{\mathbbm{k}}[\underline{x}]$ and $J\subset{\mathbbm{k}}[\underline{y}]$, where ${\mathbbm{k}}[\underline{x}]$ and ${\mathbbm{k}}[\underline{y}]$ are polynomial rings, are studied in [BHJT21, KNT24]. As an application of Theorem 2.5 and Proposition 3.3, we extend those results ideals in standard graded finitely algebra domains over algebraically closed fields.
@@
Let $I\subset A$ and $J\subset B$ be homogeneous ideals in finitely generated regular ${\mathbbm{k}}$-algebra domains $A$ and $B$, where ${\mathbbm{k}}$ is algebraically closed. Set $I+J=IR+JR$, where $R=A\otimes B$. Then the following hold .
@@ Acknowledgments.
The authors are thankful to Tài Huy Hà for suggesting the problem and valuable comments. This is a part of the second author’s Ph.D. thesis. We are also thankful to Souvik Dey and Hop D. Nguyen for some helpful discussions.
@@ 2. Symbolic Powers via flat extension
Let $\phi:A\to B$ be a flat homomorphism. We set $\operatorname{Ass}_{A}^{*}(I)=\cup_{n\in{\mathbb{N}}}\operatorname{Ass}_{A}(I^{n})$. If $I\subset A$ is a ideal then we have $\operatorname{Ass}_{B}(IB)=\cup_{{\mathfrak{p}}\in\operatorname{Ass}_{A}(I)}\operatorname{Ass}_{B}({\mathfrak{p}}B)$. Furthermore, if $\operatorname{Ass}_{B}^{*}(IB)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}_{A}^{*}(I)\}$ then we have
  1. (1)
    $\operatorname{Ass}_{B}(IB)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}_{A}(I)\}$,
  2. (2)
    $\bigcup_{{\mathfrak{q}}\in\operatorname{Ass}_{B}(I^{(n)}B)}{\mathfrak{q}}\subset\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}_{A}(I)}{\mathfrak{p}}B\text{ for all }n\geqslant 1.$
@@
$\operatorname{Ass}_{B}(IB)=\operatorname{Ass}_{B}(A/I\otimes_{A}B)=\bigcup_{\mathfrak{p}\in\operatorname{Ass}_{A}(A/I)}\operatorname{Ass}_{B}(B/\mathfrak{p}B)=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\operatorname{Ass}_{B}({\mathfrak{p}}B).$
By the assumption ${\mathfrak{p}}B$ is prime for each ${\mathfrak{p}}\in\operatorname{Ass}_{A}(I)$. Hence we get
$\operatorname{Ass}_{B}(IB)=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\operatorname{Ass}_{B}({\mathfrak{p}}B)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}(I)\}.$
Note that by [HJKN23, Lemma 2.2] if ${\mathfrak{p}}\in\operatorname{Ass}(I^{(n)})$ then ${\mathfrak{p}}\subset{\mathfrak{q}}$ for some ${\mathfrak{q}}\in\operatorname{Ass}(I)$. Thus we get $\cup_{{\mathfrak{p}}\in\operatorname{Ass}(I^{(n)})}{\mathfrak{p}}\subset\cup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}$ for all $n\geqslant 1$. Since $\operatorname{Ass}(IB)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}(I)\}$ then using Theorem 2.3 we get
$\bigcup_{{\mathfrak{q}}\in\operatorname{Ass}(I^{(n)}B)}{\mathfrak{q}}=\bigcup_{\begin{subarray}{c}{\mathfrak{q}}\in\operatorname{Ass}({\mathfrak{p}}B),\\ {\mathfrak{p}}\in\operatorname{Ass}\left(I^{(n)}\right)\end{subarray}}{\mathfrak{q}}=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I^{(n)})}{\mathfrak{p}}B\subset\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}B,\text{ as }\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}\left(I^{(n)}\right)}\subset\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}.$
@@
Let $A$ and $B$ be Noetherian rings and $\phi:A\to B$ be a flat morphism. Assume that $I\subset A$ is a proper ideal such that $\operatorname{Ass}^{*}(IB)=\{{\mathfrak{p}}B~{}|~{}\text{ for all }{\mathfrak{p}}\in\operatorname{Ass}^{*}(I)\}$, and every associated prime of $I$ is a contraction of some ideal in $B$. Then $(IB)^{(n)}=I^{(n)}B$, for all $n\geqslant 1$.
@@
First we check that as in Lemma 2.4, $\operatorname{Ass}_{B}(IB)=\{{{\mathfrak{p}}}B~{}|~{}{{\mathfrak{p}}}\in\operatorname{Ass}_{A}(I)\}$. Using the definition of symbolic powers of an ideal and assumption, we get
@@
Now by Lemma 2.4, we have $\bigcup_{\operatorname{Ass}{\mathfrak{q}}\in\left(I^{(n)}B\right)}{\mathfrak{q}}\subset\bigcup_{{\mathfrak{q}}\in\operatorname{Ass}(IB)}{\mathfrak{q}}=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}B$. Thus whenever $s\in W$ then $s\not\in{\mathfrak{q}}$, where ${\mathfrak{q}}$ is an associated prime of $I^{(n)}B$. Thus by [AM94, Proposition 4.4(iii)] we get
@@ 3. Resurgence and Prime extension
Since $\phi$ is prime extension then by remark 3.2, it is faithfully flat. Thus using Theorem 2.5 and [AM94, ex 3.16(I)] we get that
@@
We want to mention that [Wal18, Proposition 2.1] also proved the similar result.
@@ 4. Resurgence of sum of ideals
In this section we prove similar results for finitely generated regular ${\mathbbm{k}}$-algebra domains, where ${\mathbbm{k}}$ is an algebraically closed field. One of the key tools in [BHJT21, Theorem 2.6] and [KNT24, Theorem 2.3] is the binomial expansion of symbolic powers of ideals, which holds for ideals in ${\mathbbm{k}}$-algebras. The other tool is $\rho(IR)=\rho(I)$, this holds in the polynomial ring set up. To generalize [BHJT21, Theorem 2.6] and [KNT24, Theorem 2.3], we need that $\rho(IR)=\rho(I)$ in general domains, which is not known.
In Lemma 4.1, we prove that if $A$ and $B$ are finitely generated algebra domains over an algebraically closed field ${\mathbbm{k}}$, and $R=A\otimes B$, then $\rho(IR)=\rho(I)$, and $\rho(JR)=\rho(J)$, where $I\subset A,J\subset B$ are ideals. In this section, we assume that all things are regular.
@@
Now $\left(A/\mathfrak{p}\right)\otimes_{k}B$ is an integral domain by [Iit82, Lemma 1.54]. It is a well-known fact that an ideal is prime if only if the quotient ring is an integral domain, so $\mathfrak{p}R$ is prime, and thus, $\phi$ has prime extension property. The last equality follows from Proposition 3.3. ∎
@@
To see why we need the conditions on $A$ and $B$, we will need some examples and theorems. Now [Vám78, Theorem 11] implies that, if $A$ (as a field extension over ${\mathbbm{k}}$) is not finitely generated as a ${\mathbbm{k}}$-algebra, then $A\otimes_{\mathbbm{k}}A$ is not a Noetherian ring. For an example, $\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}$ is not a Notherian ring. Therefore, the condition being finitely generated is necessary. We also need $A$ and $B$ to be integral domains as ${\mathbbm{k}}$-algebras where ${\mathbbm{k}}$ is algebraically closed because of [Iit82, Lemma 1.54], this will make $A\otimes_{\mathbbm{k}}B$ to be an integral domain. It’s a famous example that $\mathbb{C}\otimes_{\mathbb{R}}\mathbb{C}$ is not an integral domain. Also note by [TY03, Theorem 6(e)] $A\otimes B$ is regular.